OpenStudy (anonymous):

Find an equation of the line through the given points. Write the equation in standard form. Through (7, -5) and (0, -3) Choose one answer. a. 12x + 3y = -9 b. 2x + 7y = -21 c. -2x + 7y = -21 d. -12x - 3y = -9

OpenStudy (mertsj):

Did you find the slope?

OpenStudy (ash2326):

Slope of a line passing through two points (7, -5) and (0, -3) is given by m=(-5-(-3))/(7-0)=-2/7 so our equation will be of the form y=-2/7 x+c for C?? put 0,-3 in the equation -3=c y=-2/7 x-3 or 7y=-2x-21 7y+2x=21 option b